OpenStudy (anonymous):

In the system shown below, what are the coordinates of the solution that lies in quadrant I? x^2+y^2=25 x+y=25 i stink at conic sections so can someone help me?

OpenStudy (amistre64):

 looks something like this to me

OpenStudy (anonymous):

hmm that looks like there are 2 intersecting points though :/

OpenStudy (amistre64):

there might be 2 points, cant guarentee that my sketch is accurate tho

OpenStudy (amistre64):

since x+y = 25 we know that y = -x+25 insert that "value" of y into the first equation and solve for x

OpenStudy (anonymous):

ok so then that would be x^2-x+25=25 correct?

OpenStudy (amistre64):

hmm x^2 + (25-x)^2 = 25 x^2 + x^2 - 50x + 625 = 25

OpenStudy (amistre64):

x^2 + x^2 - 50x + 625 = 25 2x^2 - 50x + 600 = 0 2(x^2 =25x + 300) = 0 x^2 =25x + 300 = 0

OpenStudy (amistre64):

.... -25 that is :)

OpenStudy (amistre64):

x^2 - 25x + 300 = 0 had to edit a typo ;)

OpenStudy (anonymous):

ok that makes more sense lol im like: uhhh what? when you put down x^2=25x+300=0 lol

OpenStudy (amistre64):

visually, i dont see an intersection; but you might wanna keep at the mathing to verify http://www.wolframalpha.com/input/?i=x%5E2%2By%5E2%3D25%2C+x%2By%3D25

OpenStudy (anonymous):

hmm.... maybe if i check with another open study person. @Mertsj ?

OpenStudy (mertsj):

y=25-x \[x^2+(25-x)^2=25\] \[x^2+625-50x+x^2=25\] \[2x^2-50x+600=0\] \[x^2-25x+300=0\]

OpenStudy (mertsj):

Can you solve that quadratic equation?

OpenStudy (mertsj):

So if you solve the quadratic equation I posted, you will find that the discriminant is negative which means there are no real solutions.

OpenStudy (anonymous):

well then why the heck does the question ask me to and i do quote: Write your answer in the form (a,b) without using spaces?

OpenStudy (mertsj):

And so the given line does not intersect the given circle in the first quadrant or anywhere else. Perhaps you have posted the equation of the line incorrectly.

OpenStudy (anonymous):

ah crap i did write it wrong its supposed to be x^2-y^2=25 and x+y=25 *facedesk*

OpenStudy (mertsj):

So you have been given the equation of a hyperbola, not a circle.

OpenStudy (mertsj):

So just use the solution I posted but change the signs.

OpenStudy (anonymous):

*facekeyboard* sorry about that lol

OpenStudy (anonymous):

change the sings? so instead of it being x^2-25+300 its now x^2+25+300=0?

OpenStudy (anonymous):

signs*

OpenStudy (mertsj):

\[x^2-(25-x)^2=25\] \[x^2-(625-50x+x^2)=25\] \[x^2-625+50x-x^2=25\] \[50x-625=25\] \[50x=650\] \[x=13\] \[y=25-13=12\] (13,12)

OpenStudy (anonymous):

double checked it and its right!! thanks mertsj!

OpenStudy (anonymous):

and thank you tcarrol010 for showing me that i plugged in the wrong equation lol

OpenStudy (mertsj):

yw